OpenStudy (anonymous):

how do I find the coterminal angle of -13pi/3?

OpenStudy (anonymous):

Here is an example with explanation.. Coterminal angles are angles, which share a common terminal side. There are two sides of an angle: the initial side and the terminal side. The initial side is the ray where the measurement of an angle starts. The terminal side is the ray where the measurement of an angle stops. 60° and 300° are coterminal. One way to find the coterminal angle to an angle is: The sum of the angle and its coterminal is 360°. Coterminal angles can be positive or negative. Ex: Find the coterminal angle for 30°. A+30°=360° --> A=330° If we let A be the coterminal angle, we could get two solutions. A=360-30 --> A=330° (the answer mentioned above.) A=-360-30 --> A=-390° Hope it helps you solve the problem.

OpenStudy (anonymous):

it didn't really help, sorry. I figured the co-terminal angle would be pi/3 or something of that nature then, but it's not.

OpenStudy (jdoe0001):

do you know what a negative angle is?

OpenStudy (jdoe0001):

like the difference between say, 90 degrees or -90 degrees?

OpenStudy (jdoe0001):

:) Skiller8860 for some reason your nick has a ring to a prime meatcut

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

didn't have to, is about 65 degrees here =)

OpenStudy (jdoe0001):

since I got fanned, lemme get my jacket on =)

OpenStudy (jdoe0001):

\(\boldsymbol{ -\cfrac{13\pi}{3}\implies -4\cfrac{1}{3}\implies -4\cfrac{\pi}{3}\implies -4\pi - \cfrac{\pi}{3}}\)

OpenStudy (jdoe0001):

so, you'd got around 2 revolutions, that \(\boldsymbol{4\pi}\) and then further down \(\boldsymbol{\cfrac{\pi}{3}}\) clockwise in the circumference

OpenStudy (jdoe0001):